# Efficient Representation of Gaussian Fermionic Pure States in Non-Computational Bases

### v1

@@ 5 Gaussian pure states in the $(\phi,\frac{\pi}{2},\alpha)$ basis
Unlike the previous sections the Gaussian states in the $(\phi,\frac{\pi}{2},\alpha)$ basis when $\alpha\neq 0$ does not have domain wall kind of interpretation. We conjecture that it can be obtained first by following the steps that lead the state in the $(\phi,\frac{\pi}{2},0)$ basis and then just putting $e^{-i\alpha}$ for each -- in the corresponding configuration. For example, we have
${}\left|\mathbb{R},101\right\rangle$ $=$ $\frac{1}{\sqrt{2}\mathcal{Z}^{\phi}}\Big{(}d^{\phi}_{(101)}\left|+++\right\rangle+e^{-i\alpha}c^{\phi}_{(101)}\left|++-\right\rangle-e^{-i\alpha}b^{\phi}_{(101)}\left|+-+\right\rangle$ (5.1)
-- $e^{-2i\alpha}a^{\phi}_{(101)}\left|+--\right\rangle+e^{-i\alpha}a^{\phi}_{(101)}\left|-++\right\rangle+e^{-2i\alpha}b^{\phi}_{(101)}\left|-+-\right\rangle$
-- $e^{-2i\alpha}c^{\phi}_{(101)}\left|--+\right\rangle-e^{-3i\alpha}d^{\phi}_{(101)}\left|---\right\rangle\Big{)}.$
@@ 6 Configuration probabilities
We start with the computational basis, i.e. $\sigma^{z}$ basis. It is easy to see that for the configuration $\mathcal{I}$ we will have
@@
${}P_{\mathcal{I}}=\det[\frac{\mathbb{I}-\mathbb{I}_{\mathcal{I}}.\mathbb{G}}{2}],$ (6.2)
@@
The next step is writing the probabilities for the bases in the $xy$ plane. Using the amplitude formulas of the previous sections and some simplifications for the probability of the configuration $\mathcal{S}$ we have
@@
${}\tilde{R}^{\phi}=\frac{\mathbb{H}-i\cot{\phi}\mathbb{H}.\mathbb{P}+\mathbb{P}-i\cot{\phi}\mathbb{I}}{-\mathbb{H}-i\cot{\phi}\mathbb{H}.\mathbb{P}+\mathbb{P}+i\cot{\phi}\mathbb{I}}.$ (6.4)
@@
${}P^{\phi}_{\mathcal{S}}=\frac{1}{2\mathcal{N}_{\tilde{R}^{\phi}}^{2}}|\det[\frac{(\mathbb{I}-\mathbb{I}_{\mathcal{S}}).\tilde{\mathbb{R}}^{\phi}+\mathbb{I}+\mathbb{I}_{\mathcal{S}}}{2}]|,$ (6.5)
where the diagonal matrix $\mathbb{I}_{\mathcal{S}}$ is defined for the configuration $\mathcal{S}$ by first finding the domain wall configuration and then we put $-1(+1)$ in the diagonal of the matrix $\mathbb{I}_{\mathcal{S}}$ when there is (no) domain wall at the site.
@@
${}P^{x}_{\mathcal{S}}=$ $\frac{1}{2}\det[\frac{\mathbb{I}-\mathbb{I}_{\mathcal{S}}.\mathbb{G}.\mathbb{P}}{2}],$ (6.6)
$P^{y}_{\mathcal{S}}=$ $\frac{1}{2}\det[\frac{\mathbb{P}-\mathbb{I}_{\mathcal{S}}.\mathbb{G}}{2}].$ (6.7)
@@ 6.1 Formation probabilities in the critical transverse field Ising chain
In this section, we examine the formation probabilities across all the bases we have explored so far for the ground state of the transverse field Ising chain at its critical point, considering both periodic and open boundary conditions.
Based on arguments from conformal field theory at the critical point, one would anticipate the formation probabilities to exhibit the following behavior( for those configurations that converge toward conformal boundary conditions) for the Periodic Boundary Condition (PBC) [25, 17, 16]:
@@
where $s_{\mathcal{S}}$ is a universal quantity and is dubbed as the Affleck-Ludwig boundary entropy. In the case of Open Boundary Condition (OBC), one expects a totally different behavior [26, 14]
@@
When the configuration flows to the boundary condition compatible with the natural boundary condition of the system then $h_{bcc}=0$, otherwise depending on the induced boundary condition one needs to add the corresponding conformal weight in the above formula.
@@
${}\mathbb{R}=(\mathbb{I}+\mathbb{G}).(\mathbb{I}-\mathbb{G})^{-1},$ (6.12)
@@
For OBC after considering various configurations we found that
@@
The above conclusion is consistent with what we expect from boundary CFT.
@@
independent of the value of $\phi$. For OBC the coefficient of the logarithm shows the following universal behavior:
@@
where $k$ is an integer number. For OBC we found $a=\frac{3}{8}$ when $L=6k+6$. In the case of the base configuration ++++-+ + - we were not able to find a consistent result that pinpointed the relevant system sizes. However, surprisingly in the case of the base configuration ++++++-+ + + - we found
@@ 7 Conclusions
Using our formulas, we analyzed the formation probabilities for various configurations in the ground state of the critical transverse field Ising chain, considering both periodic and open boundary conditions. This examination across different configurations revealed a coherent understanding of the universal quantities, such as boundary entropy in the periodic case and the logarithm coefficient in the open case, for all bases in the $xy$ plane except the $\sigma^{y}$ basis. In this particular basis, we noted significant anomalies for crystalline configurations with unequal numbers of up and down spins. Investigating these specific configurations through analytical methods, especially for periodic chains, could provide a comprehensive insight into the phenomena.
Looking ahead, the natural progression of this research lies in exploring Gaussian mixed states, a topic we aim to revisit in future work. Such an extension would deepen the understanding and broaden the applicability of Gaussian states in quantum systems.
@@ Acknowledgements
We thank M. Heyl for reading the manuscript and comments. MAR thanks CNPq and FAPERJ (grant number E-26/210.062/2023) for partial support.
@@ Appendix A Generic Gaussian pure states
Consider the following generic Gaussian pure state:
@@
$\begin{split}e^{\frac{1}{2}\begin{pmatrix}\bar{\mathbf{c}}^{\dagger}&\bar{\mathbf{c}}\\ \end{pmatrix}\bar{\mathbf{M}}_{\mathcal{C}}\begin{pmatrix}\tilde{\mathbf{c}}\\ \tilde{\mathbf{c}}^{\dagger}\end{pmatrix}}\left|\bar{0}\right\rangle=e^{\frac{1}{2}\bar{\mathbf{c}}^{\dagger}\bar{\mathbf{R}}\bar{\mathbf{c}}^{\dagger}}e^{\bar{\mathbf{c}}^{\dagger}\bar{\mathbf{Y}}\bar{\mathbf{c}}-\frac{1}{2}Tr\bar{\mathbf{Y}}}e^{\frac{1}{2}\bar{\mathbf{c}}\bar{\mathbf{Z}}\bar{\mathbf{c}}}\left|\bar{0}\right\rangle\to e^{\frac{1}{2}\bar{\mathbf{c}}^{\dagger}\bar{\mathbf{R}}\bar{\mathbf{c}}^{\dagger}}\left|\bar{0}\right\rangle,\end{split}$ (A.5)
@@
After writing back the state (A.5) with respect to the original creation and annihilation operators the above state has the form (2.5).
@@ Appendix B Examples of Gaussian pure states
In this Appendix, we list some examples of the Gaussian states in different bases for various sizes. We first list all the possibilities for the $L=2$ and then for $L=3$, we list them for just the $\sigma^{z}$ and $\sigma^{x}$ bases.
@@ B.1 Gaussian pure states with size $L=2$
In this section, we write the Gaussian state for $L=2$ in different bases. We first define:
@@
$\mathcal{N}_{R}$ $=$ $\sqrt{1+r_{12}^{2}},\ \mathcal{Z}^{\phi}=2\sqrt{1+r_{12}^{2}}.$
$\textbf{R}=\left(\begin{array}[]{cc}0&\text{$r_{12}$}\\ -\text{$r_{12}$}&0\\ \end{array}\right),$ (B.1)
@@
state configs
$\left|11\right\rangle$
$\left|10\right\rangle$
$\left|01\right\rangle$
$\left|00\right\rangle$
$\left|\textbf{R},00\right\rangle$
$r_{12}$
00
00
$1$
$\left|\textbf{R},01\right\rangle$
00
$r_{12}$
$1$
00
$\left|\textbf{R},10\right\rangle$
00
$1$
-$r_{12}$
00
$\left|\textbf{R},11\right\rangle$
$1$
00
00
-$r_{12}$
Table 2: The state in the $\sigma^{z}$ basis, with the state represented by a normalization prefactor of $\frac{1}{\mathcal{N}_{R}}$.
$L=2$, $\sigma^{x}$ basis
state configs
@@
[…]
$\left|\textbf{R},00\right\rangle$
$a^{\phi}_{(00)}$
-$b^{\phi}_{(00)}$
-$b^{\phi}_{(00)}$
$a^{\phi}_{(00)}$
$\left|\textbf{R},01\right\rangle$
$a^{\phi}_{(01)}$
$b^{\phi}_{(01)}$
-$b^{\phi}_{(01)}$
-$a^{\phi}_{(01)}$
$\left|\textbf{R},10\right\rangle$
$b^{\phi}_{(10)}$
-$a^{\phi}_{(10)}$
$a^{\phi}_{(10)}$
-$b^{\phi}_{(10)}$
$\left|\textbf{R},11\right\rangle$
$b^{\phi}_{(11)}$
$a^{\phi}_{(11)}$
$a^{\phi}_{(11)}$
$b^{\phi}_{(11)}$
Table 3: the state in the $\sigma^{x}$ basis, with the state represented by a normalization prefactor of $\frac{1}{\sqrt{2}\mathcal{Z}^{\phi}}$.
@@ B.2 Gaussian pure states with size $L=3$
$\bf{\tilde{R}}=\frac{1}{\text{$a$}}\left(\begin{array}[]{ccc}0&\text{$c$}&\text{$d$}\\ -\text{$c$}&0&\text{$b$}\\ -\text{$d$}&-\text{$b$}&0\\ \end{array}\right)$ (B.3)
@@
state configs
$\left|111\right\rangle$
$\left|110\right\rangle$
$\left|101\right\rangle$
$\left|100\right\rangle$
$\left|011\right\rangle$
$\left|010\right\rangle$
$\left|001\right\rangle$
$\left|000\right\rangle$
$\left|\textbf{R},000\right\rangle$
00
$r_{12}$
$r_{13}$
00
$r_{23}$
00
00
$1$
$\left|\textbf{R},001\right\rangle$
$r_{12}$
00
00
$r_{13}$
00
$r_{23}$
$1$
00
$\left|\textbf{R},010\right\rangle$
-$r_{13}$
00
00
$r_{12}$
00
$1$
-$r_{23}$
00
$\left|\textbf{R},011\right\rangle$
00
-$r_{13}$
$r_{12}$
00
$1$
00
00
-$r_{23}$
$\left|\textbf{R},100\right\rangle$
$r_{23}$
00
00
$1$
00
-$r_{12}$
-$r_{13}$
00
$\left|\textbf{R},101\right\rangle$
00
$r_{23}$
$1$
00
-$r_{12}$
00
00
-$r_{13}$
$\left|\textbf{R},110\right\rangle$
00
$1$
-$r_{23}$
00
$r_{13}$
00
00
-$r_{12}$
$\left|\textbf{R},111\right\rangle$
$1$
00
00
-$r_{23}$
00
$r_{13}$
-$r_{12}$
00
Table 4: The state in the $\sigma^{z}$ basis, with the state represented by a normalization prefactor of $\frac{1}{\mathcal{N}_{R}}$.
@@
[…]
$\left|\textbf{R},000\right\rangle$
$a$
-$b$
-$c$
$d$
-$d$
$c$
$b$
-$a$
$\left|\textbf{R},001\right\rangle$
$a$
$b$
-$c$
-$d$
-$d$
-$c$
$b$
$a$
$\left|\textbf{R},010\right\rangle$
$b$
-$a$
$d$
-$c$
-$c$
$d$
-$a$
$b$
$\left|\textbf{R},011\right\rangle$
$b$
$a$
$d$
$c$
-$c$
-$d$
-$a$
-$b$
$\left|\textbf{R},100\right\rangle$
$d$
-$c$
-$b$
$a$
$a$
-$b$
-$c$
$d$
$\left|\textbf{R},101\right\rangle$
$d$
$c$
-$b$
-$a$
$a$
$b$
-$c$
-$d$
$\left|\textbf{R},110\right\rangle$
$c$
-$d$
$a$
-$b$
$b$
-$a$
$d$
-$c$
$\left|\textbf{R},111\right\rangle$
$c$
$d$
$a$
$b$
$b$
$a$
$d$
$c$
Table 5: The state in the $\sigma^{x}$ basis, with the state represented by a normalization prefactor of $\frac{1}{\sqrt{2}\mathcal{Z}}$.
@@
In this Appendix, we list the exact form of the states $\left|\mathbb{R},101000\right\rangle$ and $\left|\mathbb{R^{\prime}},011101\right\rangle$ which is used to get the equation (2.13).
@@
-- $\mathbb{pf}\ R_{3456}\left|100111\right\rangle+\mathbb{pf}\ R\left|010111\right\rangle$
).).) . (C.1)
@@
$+$ $r_{24}\left|001001\right\rangle-r^{\prime}_{25}\left|001111\right\rangle-r^{\prime}_{26}\left|001100\right\rangle-r^{\prime}_{34}\left|010001\right\rangle+r_{35}\left|010111\right\rangle+r^{\prime}_{36}\left|010100\right\rangle$
-- $r^{\prime}_{45}\left|011011\right\rangle-r^{\prime}_{46}\left|011000\right\rangle+r_{56}\left|011110\right\rangle-\mathbb{pf}\ R^{\prime}_{1234}\left|100001\right\rangle+\mathbb{pf}\ R^{\prime}_{1235}\left|100111\right\rangle$
$+$ $\mathbb{pf}\ R_{1236}\left|100100\right\rangle-\mathbb{pf}\ R^{\prime}_{1245}\left|101011\right\rangle-\mathbb{pf}\ R^{\prime}_{1246}\left|101000\right\rangle+\mathbb{pf}\ R^{\prime}_{1256}\left|101110\right\rangle$
@@
-- $\mathbb{pf}\ R^{\prime}_{3456}\left|1010010\right\rangle-\mathbb{pf}\ R^{\prime}\left|1100010\right\rangle$
).).) . (C.2)

### v2

@@ 5 Gaussian pure states in the $(\phi,\frac{\pi}{2},\alpha)$ basis
Unlike the previous sections, the Gaussian states in the $(\phi,\frac{\pi}{2},\alpha)$ basis when $\alpha\neq 0$ does not have a domain wall kind of interpretation. We conjecture that it can be obtained first by following the steps that lead the state in the $(\phi,\frac{\pi}{2},0)$ basis and then just putting $e^{-i\alpha}$ for each -- in the corresponding configuration. For example, we have
${}\left|\mathbb{R},101\right\rangle$ $=$ $\frac{1}{\sqrt{2}\mathcal{Z}^{\phi}}\Big{(}d^{\phi}_{(101)}\left|+++\right\rangle+e^{i\alpha}c^{\phi}_{(101)}\left|++-\right\rangle$ (5.1)
-- $e^{-i\alpha}b^{\phi}_{(101)}\left|+-+\right\rangle-e^{-2i\alpha}a^{\phi}_{(101)}\left|+--\right\rangle$
$+$ $e^{-i\alpha}a^{\phi}_{(101)}\left|-++\right\rangle+e^{-2i\alpha}b^{\phi}_{(101)}\left|-+-\right\rangle$
-- $e^{-2i\alpha}c^{\phi}_{(101)}\left|--+\right\rangle$
-- $e^{-3i\alpha}d^{\phi}_{(101)}\left|---\right\rangle\Big{)}.$
@@ 6 Configuration probabilities
We start with the computational basis, i.e. $\sigma^{z}$ basis. It is easy to see that for the configuration $\mathcal{I}$, we will have
@@
${}P_{\mathcal{I}}=\det[\frac{\mathbb{I}-\mathbb{I}_{\mathcal{I}}\cdot\mathbb{G}}{2}],$ (6.2)
@@
The next step is writing the probabilities for the bases in the $xy$ plane. Using the amplitude formulas of the previous sections and some simplifications for the probability of the configuration $\mathcal{S}$, we have
@@
${}\tilde{R}^{\phi}=\Big{(}\mathbb{H}-i\cot{\phi}\mathbb{H}\cdot\mathbb{P}+\mathbb{P}-i\cot{\phi}\mathbb{I}\Big{)}\cdot\Big{(}-\mathbb{H}-i\cot{\phi}\mathbb{H}\cdot\mathbb{P}+\mathbb{P}+i\cot{\phi}\mathbb{I}\Big{)}^{-1}$ (6.4)
@@
${}P^{\phi}_{\mathcal{S}}=\frac{1}{2\mathcal{N}_{\tilde{R}^{\phi}}^{2}}|\det[\frac{(\mathbb{I}-\mathbb{I}_{\mathcal{S}})\cdot\tilde{\mathbb{R}}^{\phi}+\mathbb{I}+\mathbb{I}_{\mathcal{S}}}{2}]|,$ (6.5)
where the diagonal matrix $\mathbb{I}_{\mathcal{S}}$ is defined for the configuration $\mathcal{S}$ by first finding the domain wall configuration, and then we put $-1(+1)$ in the diagonal of the matrix $\mathbb{I}_{\mathcal{S}}$ when there is (no) domain wall at the site.
@@
${}P^{x}_{\mathcal{S}}=$ $\frac{1}{2}\det[\frac{\mathbb{I}-\mathbb{I}_{\mathcal{S}}\cdot\mathbb{G}\cdot\mathbb{P}}{2}],$ (6.6)
$P^{y}_{\mathcal{S}}=$ $\frac{1}{2}\det[\frac{\mathbb{P}-\mathbb{I}_{\mathcal{S}}\cdot\mathbb{G}}{2}].$ (6.7)
@@ 6.1 Formation probabilities in the critical transverse field Ising chain
In this section, we explore the formation probabilities for the ground state of the transverse field Ising (TFI) chain at its critical point. We consider both periodic and open boundary conditions.
For systems like the TFI chain under critical conditions, the logarithm of formation probabilities can be interpreted as the free energy in a Conformal Field Theory (CFT) setting. Specifically, when periodic boundary conditions are used, the system resembles a CFT on a cylinder, with boundaries at both ends. Under these conditions, an additional term known as boundary entropy emerges, contributing to the bulk free energy. Conversely, when open boundary conditions are applied, the system transforms into what can be visualized as a strip. This configuration imposes a natural boundary condition that directly arises from the Hamiltonian’s properties. In such cases, the boundary conditions determined by the formation probabilities might align with or deviate from the system’s natural boundary conditions. This deviation sometimes requires the use of conformal boundary changing operators. For further details and discussion, see references [25, 17, 16].
Based on the above arguments coming from CFT, one would anticipate the formation probabilities to exhibit the following behavior( for those configurations that converge toward conformal boundary conditions) for the Periodic Boundary Condition (PBC) [25, 17, 16]:
@@
where $s_{\mathcal{S}}$ is a universal quantity and is dubbed as the Affleck-Ludwig boundary entropy. It should be noted that for all systems belonging to the Ising universality class, such as the XY chain at the critical Ising line ($h=1$), we anticipate that this quantity will be universal and take one of the three values that we will list later. For detailed study in the $\sigma^{z}$ basis, see [16]. In addition, in this study, we show that the quantity is also universal with respect to the change of basis.
In the case of Open Boundary Condition (OBC), one expects a totally different behavior [26, 14]
@@
When the configuration flows to the boundary condition compatible with the natural boundary condition of the system, then $h_{bcc}=0$; otherwise, depending on the induced boundary condition, one needs to add the corresponding conformal weight in the above formula.
@@
${}\mathbb{R}=(\mathbb{I}+\mathbb{G})\cdot(\mathbb{I}-\mathbb{G})^{-1},$ (6.12)
@@
For OBC, after considering various configurations, we found that
@@
The above conclusion is consistent with what we expect from boundary CFT. In our analysis, we considered systems of sizes as large as $L\approx 1000$, focusing our data fitting primarily on sizes within the range $(400\leq L\leq 1000)$. The same system sizes were employed across all other bases and configurations. It’s worth mentioning that the choice of $L\approx 1000$ was not due to computational limitations but rather because the fit quality is already excellent at this size. Furthermore, since the calculations involve computing a series of determinants, it is feasible to extend our analysis to system sizes of several thousand using Mathematica with relative ease. The numbers presented were obtained by fitting the logarithm of formation probabilities to equations (6.8) and (6.9) using Mathematica. The outcomes of the fitting process align within less than one percent of the values that have been reported.
@@
independent of the value of $\phi$. For OBC, the coefficient of the logarithm shows the following universal behavior:
@@
where $k$ is an integer number. For OBC we found $a=\frac{3}{8}$ when $L=6k+6$. In the case of the base configuration ++++-+ + -, we were not able to find a consistent result that pinpointed the relevant system sizes. However, surprisingly in the case of the base configuration ++++++-+ + + -, we found
@@ 7 Conclusions
Using our formulas, we analyzed the formation probabilities for various configurations in the ground state of the critical transverse field Ising chain, considering both periodic and open boundary conditions. This examination across different configurations revealed a coherent understanding of the universal quantities, such as boundary entropy in the periodic case and the logarithm coefficient in the open case, for all bases in the $xy$ plane except the $\sigma^{y}$ basis. In this particular basis, we noted significant anomalies for crystalline configurations with unequal numbers of up and down spins. Investigating these specific configurations through analytical methods, especially for periodic chains, could provide a comprehensive insight into the phenomena. In the future, we plan to delve into Gaussian mixed states to further enhance our understanding and expand the practical applications of Gaussian states in quantum systems.
@@ Acknowledgements
We thank M. Heyl for reading the manuscript and comments. We thank M. A. Seifi Mirjafarlou for pointing us to a few typos. MAR thanks CNPq and FAPERJ (grant number E-26/210.062/2023) for partial support.
@@ Appendix A Generic Gaussian pure states
In this appendix, we will show that more general Gaussian states can be written in the form (2.5). Consider the following generic Gaussian pure state:
@@
$\begin{split}&e^{\frac{1}{2}\begin{pmatrix}\bar{\mathbf{c}}^{\dagger}&\bar{\mathbf{c}}\\ \end{pmatrix}\bar{\mathbf{M}}_{\mathcal{C}}\begin{pmatrix}\bar{\mathbf{c}}\\ \bar{\mathbf{c}}^{\dagger}\end{pmatrix}}\left|\bar{0}\right\rangle=\\ &e^{\frac{1}{2}\bar{\mathbf{c}}^{\dagger}\bar{\mathbf{R}}\bar{\mathbf{c}}^{\dagger}}e^{\bar{\mathbf{c}}^{\dagger}\bar{\mathbf{Y}}\bar{\mathbf{c}}-\frac{1}{2}Tr\bar{\mathbf{Y}}}e^{\frac{1}{2}\bar{\mathbf{c}}\bar{\mathbf{Z}}\bar{\mathbf{c}}}\left|\bar{0}\right\rangle\to e^{\frac{1}{2}\bar{\mathbf{c}}^{\dagger}\bar{\mathbf{R}}\bar{\mathbf{c}}^{\dagger}}\left|\bar{0}\right\rangle,\end{split}$ (A.5)
@@
After writing back the state (A.5) with respect to the original creation and annihilation operators, the above state has the form (2.5).
@@ Appendix B Examples of Gaussian pure states
In this Appendix, we list some examples of the Gaussian states in different bases for various sizes. We first list all the possibilities for the $L=2$, and then for $L=3$ we just list them for the $\sigma^{z}$ and $\sigma^{x}$ bases.
@@ B.1 Gaussian pure states with size $L=2$
In this section, we write the Gaussian states with $L=2$ in different bases. We first define:
@@
$\mathcal{N}_{R}$ $=$ $\sqrt{1+r_{12}^{2}},$
$\mathcal{Z}^{\phi}$ $=$ $2\sqrt{1+r_{12}^{2}},$
and
$\textbf{R}=\left(\begin{array}[]{cc}0&\text{$r_{12}$}\\ -\text{$r_{12}$}&0\\ \end{array}\right).$ (B.1)
The corresponding $\mathbb{R}^{\phi}$ and ~$\tilde{\mathbb{R}}^{\phi}$ are listed in the Table 1. List of the states in the $\sigma^{z}$ and $(\phi,\frac{\pi}{2},\alpha)$ bases are listed in the Tables 2 and 3.
@@
state configs $\left|11\right\rangle$ $\left|10\right\rangle$ $\left|01\right\rangle$ $\left|00\right\rangle$
$\left|\textbf{R},00\right\rangle$ $r_{12}$ 00 00 $1$
$\left|\textbf{R},01\right\rangle$ 00 $r_{12}$ $1$ 00
$\left|\textbf{R},10\right\rangle$ 00 $1$$r_{12}$ 00
$\left|\textbf{R},11\right\rangle$ $1$ 00 00 -$r_{12}$
Table 2: The state in the $\sigma^{z}$ basis, with a normalization prefactor of $\frac{1}{\mathcal{N}_{R}}$.
$L=2$, $(\phi,\frac{\pi}{2},\alpha)$ basis
@@
$\left|\textbf{R},00\right\rangle$ $a^{\phi}_{(00)}$$b^{\phi}_{(00)}$$b^{\phi}_{(00)}$ $a^{\phi}_{(00)}$
$\left|\textbf{R},01\right\rangle$ $a^{\phi}_{(01)}$ $b^{\phi}_{(01)}$$b^{\phi}_{(01)}$$a^{\phi}_{(01)}$
$\left|\textbf{R},10\right\rangle$ $b^{\phi}_{(10)}$$a^{\phi}_{(10)}$ $a^{\phi}_{(10)}$$b^{\phi}_{(10)}$
$\left|\textbf{R},11\right\rangle$ $b^{\phi}_{(11)}$ $a^{\phi}_{(11)}$ $a^{\phi}_{(11)}$ $b^{\phi}_{(11)}$
Table 3: The state in the $(\phi,\frac{\pi}{2},\alpha)$ basis, with a normalization prefactor of $\frac{1}{\sqrt{2}\mathcal{Z}^{\phi}}$.
@@ B.2 Gaussian pure states with size $L=3$
and
$\bf{\tilde{R}}=\frac{1}{\text{$a$}}\left(\begin{array}[]{ccc}0&\text{$c$}&\text{$d$}\\ -\text{$c$}&0&\text{$b$}\\ -\text{$d$}&-\text{$b$}&0\\ \end{array}\right).$ (B.3)
List of all the states in the $\sigma^{z}$ and $\sigma^{x}$ bases are in the tables 4 and 5.
@@
state configs $\left|111\right\rangle$ $\left|110\right\rangle$ $\left|101\right\rangle$ $\left|100\right\rangle$ $\left|011\right\rangle$ $\left|010\right\rangle$ $\left|001\right\rangle$ $\left|000\right\rangle$
$\left|\textbf{R},000\right\rangle$ 00 $r_{12}$ $r_{13}$ 00 $r_{23}$ 00 00 $1$
$\left|\textbf{R},001\right\rangle$ $r_{12}$ 00 00 $r_{13}$ 00 $r_{23}$ $1$ 00
$\left|\textbf{R},010\right\rangle$$r_{13}$ 00 00 $r_{12}$ 00 $1$$r_{23}$ 00
$\left|\textbf{R},011\right\rangle$ 00 -$r_{13}$ $r_{12}$ 00 $1$ 00 00 -$r_{23}$
$\left|\textbf{R},100\right\rangle$ $r_{23}$ 00 00 $1$ 00 -$r_{12}$$r_{13}$ 00
$\left|\textbf{R},101\right\rangle$ 00 $r_{23}$ $1$ 00 -$r_{12}$ 00 00 -$r_{13}$
$\left|\textbf{R},110\right\rangle$ 00 $1$$r_{23}$ 00 $r_{13}$ 00 00 -$r_{12}$
$\left|\textbf{R},111\right\rangle$ $1$ 00 00 -$r_{23}$ 00 $r_{13}$$r_{12}$ 00
Table 4: The state in the $\sigma^{z}$ basis, with a normalization prefactor of $\frac{1}{\mathcal{N}_{R}}$.
@@
$\left|\textbf{R},000\right\rangle$ $a$$b$$c$ $d$$d$ $c$ $b$$a$
$\left|\textbf{R},001\right\rangle$ $a$ $b$$c$$d$$d$$c$ $b$ $a$
$\left|\textbf{R},010\right\rangle$ $b$$a$ $d$$c$$c$ $d$$a$ $b$
$\left|\textbf{R},011\right\rangle$ $b$ $a$ $d$ $c$$c$$d$$a$$b$
$\left|\textbf{R},100\right\rangle$ $d$$c$$b$ $a$ $a$$b$$c$ $d$
$\left|\textbf{R},101\right\rangle$ $d$ $c$$b$$a$ $a$ $b$$c$$d$
$\left|\textbf{R},110\right\rangle$ $c$$d$ $a$$b$ $b$$a$ $d$$c$
$\left|\textbf{R},111\right\rangle$ $c$ $d$ $a$ $b$ $b$ $a$ $d$ $c$
Table 5: The state in the $\sigma^{x}$ basis, with a normalization prefactor of $\frac{1}{\sqrt{2}\mathcal{Z}}$.
@@
In this Appendix, we list the exact form of the states $\left|\mathbb{R},101000\right\rangle$ and $\left|\mathbb{R^{\prime}},011101\right\rangle$ which is used to get the equation (2.13). We have
@@
-- $\mathbb{pf}\ R_{3456}\left|100111\right\rangle+\mathbb{pf}\ R\left|010111\right\rangle)$
and
@@
$+$ $r^{\prime}_{24}\left|001001\right\rangle-r^{\prime}_{25}\left|001111\right\rangle-r^{\prime}_{26}\left|001100\right\rangle-r^{\prime}_{34}\left|010001\right\rangle+r^{\prime}_{35}\left|010111\right\rangle+r^{\prime}_{36}\left|010100\right\rangle$
-- $r^{\prime}_{45}\left|011011\right\rangle-r^{\prime}_{46}\left|011000\right\rangle+r^{\prime}_{56}\left|011110\right\rangle-\mathbb{pf}\ R^{\prime}_{1234}\left|100001\right\rangle+\mathbb{pf}\ R^{\prime}_{1235}\left|100111\right\rangle$
$+$ $\mathbb{pf}\ R^{\prime}_{1236}\left|100100\right\rangle-\mathbb{pf}\ R^{\prime}_{1245}\left|101011\right\rangle-\mathbb{pf}\ R^{\prime}_{1246}\left|101000\right\rangle+\mathbb{pf}\ R^{\prime}_{1256}\left|101110\right\rangle$
@@
-- $\mathbb{pf}\ R^{\prime}_{3456}\left|010010\right\rangle-\mathbb{pf}\ R^{\prime}\left|100010\right\rangle).$